Assembly

What are the contents of memory locations 200, 201, 202, and 203if
the word 1234 is stored at offset 200 and the word 5678 is storedat
offset 202?
****************************************************************************
Calculate the physical memory address generated by thefollowing
segment offset pairs.
a. 1DDD:0436
b. 1234:7920
c. 74F0:2123
d. 0000:6727
e. FFFF:4336
f. 1080:0100
g. AB01:FFFF
Answer #1
Пnhÿ6r,The physical memory address generated by the following segment offset pairs.
1a. 1DDD:0436 --------- 1 D D D 0+0 4 3 6--------------1E 2 0 6 = physical memory addressb. 1234:7920--------- 12 34 0+792 0--------------19 C6 0 = physical memory address
c. 74F0:2123 ---------74F0 0+2123--------------77023 = physical memory address
d. 0000:6727 ---------0000 0+6727--------------0672 7 = physical memory address1 1 1
e. FFFF:4336 ---------FFFF 0+433 6--------------10 4 3 2 6 = physical memory address
f. 1080:0100 --------- 1080 0+010 0--------------10 9 0 0 = physical memory address211
g. AB01:FFFF ---------A B 0 1+FF F F--------------&õ;nhÿ6Hà;21 1 00 = physical memory address
I hope this will helps you!please post the other question another post!onë'bpxk5ø<<font color="red" face="Verdana"å-e"ä.ø<>f'(0)*
answered by: Krystal